what is one of the complementary angles is three more than two that twice the other? find the measure of each?

Answer by nerdybill(5404)

what is one of the complementary angles is three more than two that twice the other? find the measure of each?
.
If two angles are complementary, the sum of the two angles equal 90 degrees.
.
Let x = one angle
then
2x+3 = other angle
.
x + 2x+3 = 90
3x+3 = 90
3x = 87
x = 29 degrees (one angle)
.
Other angle:
2x+3 = 2(29)+3 = 58+3 = 61 degrees